# Decadal analysis of sea surface temperature patterns, climatology, and anomalies in temperate coastal waters with Landsat-8 TIRS observations

### v1

@@ Abstract
Sea surface temperature (SST) is a fundamental physical parameter characterising the thermal state of sea surface. Due to the intricate thermal interactions between land, sea, and atmosphere, the spatial gradients of SST in coastal waters often appear at finer spatial scales than those in open ocean waters. The Thermal Infrared Sensor (TIRS) onboard Landsat-8, with its 100-meter spatial resolution, offers a unique opportunity to uncover fine-scale coastal SST patterns that would otherwise be overlooked by coarser-resolution thermal sensors. In this study, we first develop an operational approach for SST retrieval from the TIRS sensor, and subsequently propose a novel algorithm for establishing daily SST climatology which serves as the baseline to detect anomalous SST events. We applied the proposed methods to temperate coastal waters in South Australia for the ten-year period from 2014 to 2023. For ground validation purposes, a buoy was deployed off the coast of Port Lincoln, South Australia, to record in-situ time-series SST. The spatiotemporal patterns of SST in the study area were analysed based on the ten years of satellite-derived SST imagery. The daily baseline climatology of SST with 100 m resolution was constructed, which allowed for the detection and analysis of anomalous SST events during the study period of 2014–2023. Our results suggest the following: (1) the satellite-derived SST data, generated with the proposed algorithm, aligned well with the in-situ measured SST values; (2) the semi-enclosed, shallow regions of Upper Spencer Gulf and Upper St Vincent Gulf showed higher temperatures during summer and cooler temperatures during winter than waters closer to the open ocean, resulting in a higher seasonal variation in SST; (3) the near-shore shallow areas in Spencer Gulf and St Vincent Gulf, and regions surrounding Kangaroo Island, were identified to have a higher probability of SST anomalies compared to the rest of the study area; and (4) anomalous SST events were more likely to happen during the warm months (January–March and October–December) than the cool months (April–September). These findings may offer valuable guidance to inform the local fishing and aquaculture industries in optimising their practices. The SST product and baseline SST climatology derived in this study have been made available for public access (please refer to the Data Availability section for details).
@@
Sea surface temperature; Landsat-8 TIRS; coastal waters; radiative transfer; thermal infrared; spatiotemporal analysis; baseline climatology; anomaly detection
@@ 1 Introduction
High-quality SST retrieval from top-of-atmosphere thermal infrared measurements requires accurate correction of atmospheric effects [1, 14]. Water-emitted thermal radiance at sea level differs from that measured by spaceborne sensors due to the interference of atmosphere in the following two respects. Firstly, the thermal radiance emitted from water is scattered and absorbed by constituents of atmosphere, resulting in an attenuated signal to be captured by the sensor. Secondly, the atmosphere also emits thermal radiance by itself, superimposing water-emitted radiance. Such atmospheric interference should be appropriately corrected, in order to recover the thermal signals radiated by sea water. Considering that the circumstance of atmosphere shifts from time to time and from location to location, it would be preferred to utilise specific atmospheric conditions at the time and location of each satellite image acquisition, rather than those sampled from a standard atmosphere database, to ensure the quality of atmospheric correction.
@@
Although Landsat satellites are primarily designed for observing Earth’s landmasses, the image footprints are able to cover sea waters close to coastline, making them suitable for coastal water monitoring. In particular, the Thermal Infrared Sensor (TIRS) onboard NASA/USGS Landsat-8 has been deemed as adequate for quality retrieval of SST [14, 15, 16, 17, 18]. The instrument measures thermal radiance at two separate spectral channels, i.e., 10.60–11.19 µm (Band 10) and 11.50–12.51 µm (Band 11) [19]. Previous studies have reported that single-channel algorithms [20, 21] based on rigorous radiative transfer modelling demonstrated adequate accuracy in surface temperature retrieval over both land [22] and water [14]. The TIRS has a spatial resolution of 100 m [19], finer than the thermal bands of earlier Landsat satellites and several other spaceborne instruments like MODIS (1 km) [23], VIIRS (750 m) [24], and Sentinel-3 SLSTR (1 km) [25]. Given that the spatial gradients of SST in coastal waters often present at smaller spatial scales than those in open ocean waters, high-resolution SST data such as those derived from TIRS may help enhance our understanding of coastal SST patterns.
@@
The temperate coastal waters of South Australia, including Spencer Gulf, St Vincent Gulf, and their adjacent waters, are ecologically significant and economically vital [26]. This region supports several local industries, including fishing and aquaculture, and is home to a range of marine species, such as the endangered Great White Shark [27]. Informed decision-making for local communities and industries depends on high-quality SST climatology data specific to South Australia’s coastal waters. The SSTAARS (Sea Surface Temperature Atlas of the Australian Regional Seas) dataset [28] has been a reliable source of baseline SST climatology, supporting studies on marine heatwave detection (e.g., [29, 30]) and ocean current analysis (e.g., [31, 32]). However, its spatial resolution, at a relatively coarse 2 km, may limit its applicability in studies requiring finer-scale oceanographic details. In particular, the complex meandering coastline and small-scale SST features in South Australia’s coastal waters necessitate a finer spatial resolution for SST climatology to capture the localised thermal dynamics. The 100 m resolution TIRS observations thus present an opportunity to extract SST climatology with enhanced spatial detail compared to state-of-the-art climatology products such as SSTAARS.
@@
In this study, we aim to analyse the SST patterns, climatology, and anomalies in temperate coastal waters off the South Australian coast with high-resolution satellite observations for the ten-year period from 2014 to 2023. We first develop an operational approach for SST retrieval from the Landsat-8 TIRS sensor, followed by proposing a novel algorithm for establishing daily SST climatology which serves as the baseline for detecting anomalous SST events. Specifically, this work aims to achieve the following objectives.
@@
    We develop an operational approach for high-quality SST retrieval from the Landsat-8 TIRS sensor, with rigorous radiative transfer modelling and spatiotemporally specific atmospheric data being applied for correcting atmospheric effects;
@@
    We propose a novel algorithm to establish the daily baseline SST climatology at a spatial resolution of 100 m, which is finer than state-of-the-art climatology datasets such as the 2 km resolution SSTAARS [28];
@@
    We compile the probability map of anomalous SST events in the study area with SST images over 2014–2023, and analyse the spatial and seasonal distribution of these anomalies; and
@@
    We introduce a tile-based parallelisation scheme for efficient processing across large spatial scales, with the derived SST product and baseline SST climatology for the study area being made available for public access (please refer to the Data Availability section for details on data access).
@@ 2.1 Study Area
This work was focused on Australian territorial waters off the coast of South Australia (134°30’00”–140°00’00”E, 32°25’58”–36°30’00”S), including Spencer Gulf, St Vincent Gulf, and their adjacent waters, as shown in Figure 1. The shape files defining the coastline and territorial water limit are available from the Australian Bureau of Statistics [37]. Major townships surrounding these waters include Adelaide (the capital city of South Australia, located on the east bank of St Vincent Gulf), Port Lincoln (located on the west bank of Spencer Gulf), and Port Augusta (situated at the north end of Spencer Gulf) (Figure 1). The Kangaroo Island is located to the south of Spencer Gulf and St Vincent Gulf (Figure 1). This temperate coastal region has a Mediterranean climate, with hot and dry summers, and mild and wet winters. Air temperatures range from 18.5°C to 29.2°C in January, and from 7.2°C to 16.9°C in July. The area receives moderate rainfall, averaging 277.5 mm per year, with most of it falling during the winter months, according to statistics from Australian Bureau of Meteorology [^1] http://www.bom.gov.au/climate/averages/tables/cw_018103.shtml.
@@ 2.2 Satellite Imagery
As the thermal infrared instrument onboard NASA/USGS Landsat-8 satellite, TIRS is capable of measuring thermal radiation emitted from Earth’s surface. It records thermal radiance at two separate spectral channels, i.e., 10.60–11.19 µm (Band 10) and 11.50–12.51 µm (Band 11), with a spatial resolution of 100 m. The TIRS has a swath width of 185 kilometres, aligning with the coverage of Operational Land Imager (OLI) onboard the same satellite. In this study, the Landsat-8 Collection 2 Level-1C TIRS product was used. It consists of top-of-atmosphere thermal infrared data with radiometrical calibration and geometrical correction having been applied. TIRS images covering the study area from 2014 to 2023 were queried within the AquaWatch Data Service (ADS) built upon CSIRO’s Earth Analytics and Science Innovation (EASI) platform. An Open Data Cube (ODC) environment [38, 39] was set up on the platform to facilitate data processing. The total number of satellite observations from 2014 to 2023 varies across the study area, ranging from 127 to 431 records, as shown in Figure 2a. Regions with overlapping satellite scenes show a higher observation frequency. Cloud cover is frequent in the study region, as shown in Figure 2b, with the percentage of cloudy observations ranging from 30.6% to 72.9%. The lowest cloud percentage is observed in the Upper Spencer Gulf, and the highest over the ocean south of Kangaroo Island. In this study, pixels flagged as cloud or cirrus by the Landsat Collection 2 Pixel Quality Assessment Band were masked out from our processing. In addition to these Landsat-8 TIRS images, we also queried Landsat-8 OLI Level-2 surface reflectance images and MODIS Terra Level-2 SST images for quality control and intercomparison purposes.
@@ 2.3 Coincident Atmosphere Data
The atmosphere data spatiotemporally coincident with each satellite observation were retrieved from the CAMS global reanalysis (EAC4) dataset [40], provided by the Copernicus Atmosphere Monitoring Service (CAMS) at the European Centre for Medium-Range Weather Forecasts (ECMWF). We chose to use spatiotemporally specific atmosphere data at the time and location of each satellite observation, rather than the data of a standard atmosphere, aiming to improve the modelling accuracy of SST retrievals. The EAC4 dataset provides high-quality atmosphere data at a 3-hour temporal interval and a spatial resolution of approximately 80 km. These data were queried using the Climate Data Store Application Program Interface (CDSAPI) from the Atmosphere Data Store, where the data were provided in an interpolated 0.75°×0.75° lon/lat grid. The vertical profiles of air pressure (in unit of Pa; Figure 3a), air temperature (in K; Figure 3b), specific humidity (in kg/kg; Figure 3c), ozone content (in kg/kg; Figure 3d), and nitrogen dioxide content (in kg/kg; Figure 3e) were retrieved. These profiles were provided at 60 vertical levels of the standard atmosphere model as defined by the United States Standard Atmosphere, 1976 [41]. It is worth noting that the model levels are associated with air pressure rather than elevation. We therefore interpreted these model levels according to [42], and converted them into altitudes (in km relative to the sea level). The vertical profiles of non-trace gasses, including nitrogen, oxygen, and carbon dioxide contents (all in kg/kg; Figures 3f–h), were calculated based on their percentage constitutions in dry air moderated by the vertical distribution of specific humidity. The vertical profiles of water vapour pressure and dry air pressure (both in Pa; Figures 3i–j) were derived from air pressure and specific humidity based on the physical equations documented in [43, 44]. The vertical profile of air density (in kg/m3; Figure 3k) was then derived from dry air pressure and water vapour pressure based on the Ideal Gas Law. Finally, the vertical profile of water vapour content (in kg/m3; Figure 3l) was converted from air density and specific humidity. The detailed procedure of deriving these atmospheric profiles is described later in Subsection 3.4. These atmosphere data were applied to SST modelling in this study.
@@ 2.4 Ground Truth Measurements
To validate the satellite-derived SST data in this study, we deployed a buoy off the coast of Port Lincoln, South Australia (135°54’01”E, 34°43’03”S), with support from CSIRO’s AquaWatch Mission, as shown in Figure 4. A YSI EXO2 multiparameter sonde is mounted onto the buoy which streams in-situ temperature measurements every 10 min at a nominal depth of 1 m. The in-situ sensor has been in operation since 31 August 2022, with regular services and calibrations being conducted to ensure the quality of the data. These in-situ temperature measurements served as a reference for validating the satellite-derived SST data. For each satellite observation, the derived SST pixels covering a buffer zone of 3 km radius (excluding land pixels and near-shore pixels within 300 m to the coastline; Figure 4b), centred at the buoy location, were averaged and compared to the in-situ temperature reading closest to the acquisition time of satellite imagery. This averaging process aimed to increase the quantity of data points and provide a more robust SST estimate.
@@
As an extended experiment to evaluate the performance of the proposed SST retrieval algorithm for temperate coastal waters beyond the study area, we sourced in-situ SST measurements from three buoys off the southeast coast of Australia [45]. These buoys, managed by the Manly Hydraulics Laboratory, provide high-quality SST records in New South Wales coastal waters. The SST records were collected using a thermistor located at the base of the buoy approximately 0.5 meters below the water surface [45]. Table 2.4 details the station name, station code, location, and record period for each buoy.

### v2

@@ Abstract
Sea surface temperature (SST) is a fundamental physical parameter characterising the thermal state of sea surface. Due to the intricate thermal interactions between land, sea, and atmosphere, the spatial gradients of SST in coastal waters often appear at finer spatial scales than those in open ocean waters. The Thermal Infrared Sensor (TIRS) onboard Landsat-8, with its 100-meter spatial resolution, offers a unique opportunity to uncover fine-scale coastal SST patterns that would otherwise be overlooked by coarser-resolution thermal sensors. In this study, we first analysed the spatiotemporal patterns of SST in South Australia’s temperate coastal waters from 2014 to 2023 by developing an operational approach for SST retrieval from the Landsat-8 TIRS sensor. A buoy was deployed off the coast of Port Lincoln, South Australia, to validate the quality of SST retrievals. Then the daily baseline climatology of SST with 100 m resolution was constructed, which allowed for the detection and analysis of anomalous SST events. Our results suggest the following: (1) the satellite-derived SST data aligned well with the in-situ measured SST values; (2) the semi-enclosed, shallow regions of Upper Spencer Gulf and Upper St Vincent Gulf showed higher temperatures during summer and cooler temperatures during winter than waters closer to the open ocean, resulting in a higher seasonal variation in SST; (3) the near-shore shallow areas in Spencer Gulf and St Vincent Gulf, and regions surrounding Kangaroo Island, were identified to have a higher probability of SST anomalies compared to the rest of the study area; and (4) anomalous SST events were more likely to happen during the warm months than the cool months. We hope these findings would be helpful in supporting the fishing and aquaculture industries in the coastal waters of South Australia. The SST product and baseline SST climatology derived in this study have been made available for public access (please refer to the Data Availability section for details).
@@
Sea surface temperature; thermal infrared; spatiotemporal analysis; baseline climatology; anomaly detection
@@ 1 Introduction
High-quality SST retrieval from top-of-atmosphere thermal infrared measurements requires accurate correction of atmospheric effects [1, 14]. Water-emitted thermal radiance at sea level differs from that measured by spaceborne sensors due to the interference of atmosphere [1], which should be appropriately corrected, in order to recover the thermal signals radiated by sea water [14]. Considering that the circumstance of atmosphere shifts from time to time and from location to location, it would be preferred to utilise specific atmospheric conditions at the time and location of each satellite image acquisition, rather than those sampled from a standard atmosphere database, to ensure the quality of atmospheric correction.
@@
Although Landsat satellites are primarily designed for observing Earth’s landmasses, the image footprints are able to cover sea waters close to coastline, making them suitable for coastal water monitoring. In particular, the Thermal Infrared Sensor (TIRS) onboard NASA/USGS Landsat-8 has been deemed as adequate for quality retrieval of SST [14, 15, 16, 17, 18]. The instrument measures thermal radiance at two separate spectral channels, i.e., 10.60–11.19 µm (Band 10) and 11.50–12.51 µm (Band 11) [19]. Despite that the two-channel design allows a split-window approach to be applied, previous studies have reported that single-channel algorithms [20, 21] based on rigorous radiative transfer modelling demonstrated superior accuracy in surface temperature retrieval from TIRS observations over both land (e.g., [22]) and water (e.g., [14]). The TIRS has a spatial resolution of 100 m [19], finer than the thermal bands of earlier Landsat satellites and several other spaceborne instruments like MODIS (1 km) [23], VIIRS (750 m) [24], and Sentinel-3 SLSTR (1 km) [25]. Given that the spatial gradients of SST in coastal waters often present at smaller spatial scales than those in open ocean waters, high-resolution SST data such as those derived from TIRS may help enhance our understanding of coastal SST patterns.
@@
The temperate coastal waters of South Australia, including Spencer Gulf, St Vincent Gulf, and their adjacent waters, are ecologically significant and economically vital [26]. This region supports several local industries, including fishing and aquaculture, and is home to a range of marine species, such as the endangered Great White Shark [27]. Informed decision-making for local communities and industries depends on high-quality SST climatology data specific to South Australia’s coastal waters. The SSTAARS (Sea Surface Temperature Atlas of the Australian Regional Seas) dataset [28] has been a reliable source of baseline SST climatology (e.g., [29, 30, 31, 32]), but its spatial resolution, at a relatively coarse 2 km, may limit its applicability in studies requiring finer-scale oceanographic details. In particular, the complex meandering coastline and small-scale SST features in South Australia’s coastal waters necessitate a finer spatial resolution for SST climatology to capture the localised thermal dynamics. Previous studies reported that, fine-scale thermal gradients at submesoscale resolution (0.1–10 km), together with larger thermal features such as mesoscale jets and eddies (10–100 km), form key drivers of marine ecosystem response in coastal waters [33, 34]. The 100 m resolution TIRS observations thus present an opportunity to extract SST climatology with enhanced spatial detail that would support the analysis of ecological processes and inform aquacultural decisions.
@@
In this study, we focus on analysing the SST patterns, climatology, and anomalies in temperate coastal waters off the South Australian coast with high-resolution satellite observations for the ten-year period from 2014 to 2023. We develop several operational approaches for high-resolution SST monitoring in South Australian coastal waters, including:
@@
    Implementing an operational approach for high-quality SST retrieval from the Landsat-8 TIRS sensor, with rigorous radiative transfer modelling and spatiotemporally specific atmospheric data being applied for correcting atmospheric effects;
@@
    Establishing the daily baseline SST climatology at a spatial resolution of 100 m for South Australian coastal waters;
@@
    Compiling the probability map of anomalous SST events in the study area with SST images over 2014–2023, and analyse the spatial and seasonal distribution of these anomalies; and
@@
    Implementing a tile-based parallelisation scheme for efficient processing across large spatial scales, with the derived SST product and baseline SST climatology for the study area being made available for public access (please refer to the Data Availability section for details on data access).
Through quantitative analysis of the data generated from the above developments, we sought to answer primarily the following questions: (1) What are the spatial and seasonal patterns of SST in South Australian coastal waters? (2) What are the characteristics of SST climatology in the study area? (3) Which regions within the study area experienced more frequent anomalous SST events during 2014–2023?
@@ 2.1 Study Area
This work was focused on Australian territorial waters off the coast of South Australia (134°30’00”–140°00’00”E, 32°25’58”–36°30’00”S), including Spencer Gulf, St Vincent Gulf, and their adjacent waters, as shown in Figure 1. The South Australian territorial waters were chosen as the study area because they are ecologically significant as habitat for species such as the endangered Great White Shark, and economically vital due to their role in supporting commercial fisheries and aquaculture industries. This coastal region falls within the spatial coverage of the Landsat-8 TIRS sensor. The shapefiles defining the coastline and territorial water limit were obtained from the Australian Bureau of Statistics [^1] https://www.abs.gov.au/statistics/standards/australian-statistical-geography-standard-asgs-edition-3/jul2021-jun2026/access-and-downloads/digital-boundary-files. These shapefiles contain high-resolution vector data, which were rasterised to a 100-m grid to ensure consistency with the resolution of TIRS satellite imagery. Major townships surrounding the study area include Adelaide (the capital city of South Australia, located on the east bank of St Vincent Gulf), Port Lincoln (located on the west bank of Spencer Gulf), and Port Augusta (situated at the north end of Spencer Gulf) (Figure 1). The Kangaroo Island is located to the south of Spencer Gulf and St Vincent Gulf (Figure 1). This temperate coastal region has a Mediterranean climate, with hot and dry summers, and mild and wet winters. Air temperatures range from 18.5°C to 29.2°C in January, and from 7.2°C to 16.9°C in July. The area receives moderate rainfall, averaging 277.5 mm per year, with most of it falling during the winter months, according to statistics from Australian Bureau of Meteorology [^2] http://www.bom.gov.au/climate/averages/tables/cw_018103.shtml.
@@ 2.2 Satellite Imagery
As the thermal infrared instrument onboard NASA/USGS Landsat-8 satellite, TIRS is capable of measuring thermal radiation emitted from Earth’s surface. It records thermal radiance at two separate spectral channels, i.e., 10.60–11.19 µm (Band 10) and 11.50–12.51 µm (Band 11), with a spatial resolution of 100 m. The TIRS has a swath width of 185 kilometres, aligning with the coverage of Operational Land Imager (OLI) onboard the same satellite. In this study, the Landsat-8 Collection 2 Level-1C TIRS product was used. It consists of top-of-atmosphere thermal infrared data with radiometrical calibration and geometrical correction having been applied. TIRS images covering the study area from 2014 to 2023 were queried within the AquaWatch Data Service (ADS) built upon CSIRO’s Earth Analytics and Science Innovation (EASI) platform, where an Open Data Cube (ODC) environment [39, 40] was set up to facilitate data processing. The study period of 2014–2023 was selected because 2014 marks the first full year of operational-quality data from the TIRS sensor onboard Landsat-8 following its launch in 2013, while 2023 represents the most recent year for which the EAC4 dataset was available at the time of this study. The total number of satellite observations from 2014 to 2023 varies across the study area, ranging from 127 to 431 records, as shown in Figure 2a. Regions with overlapping satellite scenes show a higher observation frequency. Cloud cover is frequent in the study region, as shown in Figure 2b, with the percentage of cloudy observations ranging from 30.6% to 72.9%. The lowest cloud percentage is observed in the Upper Spencer Gulf, and the highest over the ocean south of Kangaroo Island. In this study, pixels flagged as cloud or cirrus by the Landsat Collection 2 Pixel Quality Assessment Band were masked out from our processing. In addition to these Landsat-8 TIRS images, we also queried Landsat-8 OLI Level-2 surface reflectance images and MODIS Terra Level-2 SST images for quality control and intercomparison purposes.
@@ 2.3 Coincident Atmosphere Data
The atmosphere data spatiotemporally coincident with each satellite observation were retrieved from the CAMS global reanalysis (EAC4) dataset [41], provided by the Copernicus Atmosphere Monitoring Service (CAMS) at the European Centre for Medium-Range Weather Forecasts (ECMWF). We chose to use spatiotemporally specific atmosphere data at the time and location of each satellite observation, rather than the data of a standard atmosphere, aiming to improve the modelling accuracy of SST retrievals. The EAC4 dataset provides high-quality atmosphere data at a 3-hour temporal interval and a spatial resolution of approximately 80 km. These data were queried using the Climate Data Store Application Program Interface (CDSAPI) from the Atmosphere Data Store, where the data were provided in an interpolated 0.75°×0.75° lon/lat grid. The vertical profiles of air pressure (in unit of Pa; Figure 3a), air temperature (in K; Figure 3b), specific humidity (in kg/kg; Figure 3c), ozone content (in kg/kg; Figure 3d), and nitrogen dioxide content (in kg/kg; Figure 3e) were retrieved. These profiles were provided at 60 vertical levels of the standard atmosphere model as defined by the United States Standard Atmosphere, 1976 [42]. It is worth noting that the model levels are associated with air pressure rather than elevation. We therefore interpreted these model levels according to [43], and converted them into altitudes (in km relative to the sea level). The vertical profiles of non-trace gasses, including nitrogen, oxygen, and carbon dioxide contents (all in kg/kg; Figures 3f–h), were calculated based on their percentage constitutions in dry air moderated by the vertical distribution of specific humidity. The vertical profiles of water vapour pressure and dry air pressure (both in Pa; Figures 3i–j) were derived from air pressure and specific humidity based on the physical equations documented in [44, 45]. The vertical profile of air density (in kg/m3; Figure 3k) was then derived from dry air pressure and water vapour pressure based on the Ideal Gas Law. Finally, the vertical profile of water vapour content (in kg/m3; Figure 3l) was converted from air density and specific humidity. These atmosphere data were applied to SST modelling in this study.
@@ 2.4 Ground Truth Measurements
To validate the satellite-derived SST data in this study, we deployed a buoy off the coast of Port Lincoln, South Australia (135°54’01”E, 34°43’03”S), with support from CSIRO’s AquaWatch Mission, as shown in Figure 4. A YSI EXO2 multiparameter sonde is mounted onto the buoy which streams in-situ bulk temperature every 10 min at a nominal depth of 1 m. The in-situ sensor has been in operation since 31 August 2022, with regular services and calibrations being conducted to ensure the quality of the data. These in-situ temperature measurements served as a reference for validating the satellite-derived SST data. For each satellite observation, the derived SST pixels covering a buffer zone of 3 km radius (excluding land pixels and near-shore pixels within 300 m to the coastline; Figure 4b), centred at the buoy location, were averaged and compared to the in-situ temperature reading closest to the acquisition time of satellite imagery. This averaging process aimed to increase the quantity of data points and provide a more robust SST estimate.
@@
As an extended experiment to evaluate the performance of the proposed SST retrieval algorithm for temperate coastal waters beyond the study area, we sourced in-situ SST measurements from three buoys off the southeast coast of Australia [46]. These buoys, managed by the Manly Hydraulics Laboratory, provide high-quality SST records in New South Wales coastal waters. The records of bulk temperature were collected using a thermistor located at the base of the buoy approximately 0.5 meters below the water surface [46]. Table 2.4 details the station name, station code, location, and record period for each buoy.